How much force is repelled by two pieces of copper, each weighing 1 g, located at a distance of 100 km, if all the atoms of one and the other piece are "taken away" by electrons, ie all atoms in the samples are completely ionized ?, 




1
Expert's answer
2022-01-05T11:37:39-0500

Determine how many atoms of copper there are in 1 g of copper:


"N=nN_A=\dfrac m{\Mu}N_A."


The charge of each atom, therefore, is


"q=eZN=e\dfrac m{\Mu}ZN_A."


The force between the charges:


"F=\dfrac 1{4\pi\epsilon_0}·\dfrac{q^2}{r^2}=\dfrac 1{4\pi\epsilon_0}·\dfrac{(emZN_A)^2}{(\Mu r)^2},\\\space\\
F=1.718·10^9\text{ N}."